A pilot can fly a plane at 150 mph in calm air. A recent trip of 400 mi flying with the wind and 400 mi returning against the wind took 6 h. Find the rate of the wind.
Answer by mananth(16946)

A pilot can fly a plane at 150 mph in calm air. A recent trip of 400 mi flying with the wind and 400 mi returning against the wind took 6 h. Find the rate of the wind.
...
let speed of wind be x mph
speed of plane = 150 mph
..
against wind speed = 150-x
with wind speed = 150 +x
..
distance = 400 miles.
time with wind + time against wind = 6 hours
..
t = d/r
400/(150-x)+400/(150+x) =6
400(150+x)+400(150-x)/(150-x)(150+x) = 6
60000+400x+60000-400x= 6(150^2-x^2)
120,000=22500-6x^2
6x^2= 135000-120000
6x^2=15000
x^2=2500
x= 50 mph rate of wind
m.ananth@hotmail.ca